The Son is three years older than the daughter.
In 17 years the son will be twice as old as the daughter is now.
What is the daughters present age?
Can you please tell me all the ways this can be written out and solved?

Answer by checkley71(8403)

s+3=d & s+17=2d
s=d-3 substituting in the second equation we get d-3+17=2d or 14=d
s=13-3 s=11 or 11+17=2X14 or 28=28